OpenStudy (anonymous):

For x = 8, x + 2 = 10. This statement written in if-then form is A. If x + 2 = 10 then x = 8. B. If x = 8 then x + 2 = 10.

jimthompson5910 (jim_thompson5910):

The answer is A. If x + 2 = 10 then x = 8. because choice B could lead x = 8 to an infinite number of possible choices for equations.

jimthompson5910 (jim_thompson5910):

Basically the fact that x = 8 doesn't uniquely tell us that x +2 = 10 since we could also have x - 3 = 5 as well

Directrix (directrix):

B. If x = 8 then x + 2 = 10. I'm thinking that x = 8 is the given so it goes in the hypothesis slot (p) and that x = 2 = 10 is the conclusion so it goes in that slot (q) for p--> q.

jimthompson5910 (jim_thompson5910):

Directrix, if x = 8, then we can easily say that x + 12 = 20 for example. There's nothing unique about x+2=10